# Orbital magnetoelectric coupling of three dimensional Chern insulators

### v2

@@ Abstract
Orbital magnetoelectric effect is closely related to the band topology of bulk crystalline insulators. Typical examples include the half quantized Chern-Simons orbital magnetoelectric coupling in three dimensional (3D) axion insulators and topological insulators, which are the hallmarks of their nontrivial bulk band topology. While the Chern-Simons coupling is well defined only for insulators with zero Chern number, the orbital magnetoelectric effects in 3D Chern insulators with nonzero (layer) Chern numbers are still open questions. In this work, we propose a never-mentioned quantization rule for the layer-resolved orbital magnetoelectric response in 3D Chern insulators, the gradient of which is exactly quantized in unit of $e^{2}/h$. By theoretical analysis and numerical simulations, we demonstrate that the quantized orbital magnetoelectric response remains robust for various types of interlayer hoppings and stackings, even against disorder and lack of symmetries. We argue that the robustness has a topological origin and protected by layer Chern number. It is thus promising to observe the proposed quantized orbital magnetoelectric response in a slab of 3D Chern insulator thanks to recent experimental developments.
@@
Band insulators can be classified into different topological phases according to different topological invariants [1, 2, 3, 4]. One important and basic topological invariant is the Chern number for two dimensional (2D) systems [2, 5, 1], which describes the winding number of hybrid Wannier centers within 2D Brillouin zone [6, 7].Those time-reversal-broken band insulators with nonzero Chern number are called Chern insulators, which would exhibit quantum anomalous Hall effect [8, 9, 10]. The notion of 2D Chern insulators can be generalized to its 3D counterpart by stacking an infinite number of 2D Chern insulators along the vertical axis [11, 12, 13, 14, 15]. In principle, one can stack 2D Chern insulators along any direction so that one can define three independent Chern numbers $(C_{x},C_{y},C_{z})$, called Chern vector [16, 17], to characterize 3D band insulators. If the interlayer hopping is weak enough such that the bulk gap remains opened, such a 3D band insulator, named 3D Chern insulator, is adiabatically connected to an infinite number of decoupled 2D Chern-insulator layers.
@@
Topological 3D band insulators are known to exhibit non-trivial orbital magnetoelectric (OME) responses. The OME coupling $\alpha_{\mu\nu}$ is defined as the change in bulk electric polarization $P_{\mu}$ while applying magnetic field $B_{\nu}$, or equivalently the variation of bulk orbital magnetization $M_{\nu}$ induced by external electric field $\mathcal{E}_{\mu}$: $\alpha_{\mu\nu}=\partial P_{\mu}/\partial B_{\nu}=\partial M_{\nu}/\partial\mathcal{E}_{\mu}$ with $\mu,\nu=x,y,z$. This response is odd under inversion and time-reversal symmetry so that it vanishes if either of two symmetries is present. 3D time-reversal-protected topological insulators are classified by $\mathbb{Z}_{2}$ invariant [4, 18, 19, 20], which exhibits a half quantized bulk diagonal Chern-Simons OME response $\alpha_{\mu\nu}=e^{2}/2h\,\delta_{\mu\nu}$, modulo an integer multiple of $e^{2}/h$, once the surface state is gapped out [21, 22]. However, the Chern-Simons coupling is well defined only for bulk band insulators with zero Chern number within any 2D plane in momentum space, namely with vanishing Chern vector. This is because the expression of the Chern-Simons OME coupling coefficient involves an integration of Chern-Simons 3-form defined by the occupied Bloch states over the 3D Brillouin zone [22, 21, 23, 24, 25], which requires the existence of a smooth and periodic gauge throughout the 3D Brillouin zone [26, 27, 28]. However, such a gauge is non-existing for 3D Chern insulators due to topological obstruction.
@@
Nevertheless, nothing forbids 3D Chern insulators to have a topological OME response. Intuitively, let us consider a 3D Chern insulator with Chern vector $\textbf{C}=(0,0,C_{z})$, where $C_{z}=C(C\in\mathbb{Z},C\neq 0)$ with interlayer distance $d$. On the one hand, it is straightforward to see that the quantized anomalous Hall current in response to weak in-plane electric field would induce to an off-diagonal OME response with its coefficient varying linearly in reap space (see Sec. VII in Supplemental Material). On the other hand, if an out-of-plane uniform electric field ($\mathcal{E}_{z}$) is applied, the surface-state occupation would be different for every layer due to the vertical electrostatic potential drop. Then, the uneven distribution of electrons among the chiral edge states of each weakly-coupled Chern-insulator layer would cause a layer-resolved orbital magnetization ($M_{z}$), which may be manifested as a new type of diagonal OME effect unique to 3D Chern insulators, as schematically shown in Fig. 1(a).
Figure 1: (a) Schematic for quasi-3D Chern insulator (see text). (b) Honeycomb lattice for 2D Haldane model on which the hopping $t_{1}$ and $t_{2}\exp(i\phi)$ are marked. (c) Two ways of stacking: $AA$-stacking and $AB$-stacking. Unless mentioned ad hoc, we always use $t_{1}=1$, $t_{2}=1/3$, $\phi=\pi/3$, $t_{3}=0.2$, $t_{4}=0.15$ and the staggered mass term is zero.
Layer-resolved orbital magnetization–
@@
Although the total diagonal $\alpha_{zz}$, only relevant in 3D case, is vanishing in the presence of mirror or inversion symmetry, the layer-resolved $\alpha_{zz}^{(l)}$ (with layer index $l$), which measures the change of layer-resolved orbital magnetization $M_{z}^{(l)}$ under a weak $\mathcal{E}_{z}$, may be finite. To define $\alpha_{zz}^{(l)}$ in the thermodynamic limit, we need to define a layer-projected orbital magnetization $M_{z}^{(l)}$ using [29, 30]
$M_{z}=\frac{e}{\hbar A}{\rm Im\,Tr}\{\hat{P}x\hat{Q}(\hat{H}-\mu)\hat{Q}y\hat{P}-\hat{Q}x\hat{P}(\hat{H}-\mu)\hat{P}y\hat{Q}\},$ (1)
@@
where $\hat{P}=\sum_{E_{n}<\mu}|\psi_{n}\rangle\langle\psi_{n}|$ and $\hat{Q}=1-\hat{P}$ are the projection operators onto the occupied and unoccupied subspaces, respectively. The chemical potential $\mu$ is at charge neutrality, $x,y$ are 2D position operators and $A$ is the system area in the $x$-$y$ plane. “Tr” in the above equation means taking the trace over the whole Hilbert space. Since $\mathcal{E}_{z}$ would break the translational symmetry in the $z$ direction, it is natural to consider a slab geometry so that $M_{z}^{(l)}$ can be suitably expressed in the basis of the Bloch function
@@
where $s$ means sublattice index in the unit cell, $l$ layer index, $n$ band index and $N_{\rm L}$ number of 2D layers in the $z$ direction. Then, we can express $M_{z}$ in the Bloch function basis as
@@
$M_{z}$ $=\frac{e}{2\hbar A}{\rm Im}\big{\{}\epsilon_{z\alpha\beta}\sum_{s\,l}\sum_{\textbf{k}}\sum_{\begin{subarray}{c}n,n^{\prime}\in{\rm o},\\ m\in{\rm e}\end{subarray}}C_{sl,n}(\textbf{k})C_{sl,n^{\prime}}^{*}(\textbf{k})$ (3)
×$\times(E_{m\textbf{k}}-\mu)\langle u_{n\textbf{k}}|\partial_{k_{\alpha}}u_{m\textbf{k}}\rangle\langle\partial_{k_{\beta}}u_{m\textbf{k}}|u_{n^{\prime}\textbf{k}}\rangle\big{\}}$
@@
$-\frac{e}{2\hbar A}{\rm Im}\big{\{}\epsilon_{z\alpha\beta}\sum_{s\,l}\sum_{\textbf{k}}\sum_{\begin{subarray}{c}m,m^{\prime}\in{\rm e},\\ n\in{\rm o}\end{subarray}}C_{sl,m}(\textbf{k})C_{sl,m^{\prime}}^{*}(\textbf{k})$
@@
×$\times(E_{n\textbf{k}}-\mu)\langle u_{m\textbf{k}}|\partial_{k_{\alpha}}u_{n\textbf{k}}\rangle\langle\partial_{k_{\beta}}u_{n\textbf{k}}|u_{m^{\prime}\textbf{k}}\rangle\big{\}}$
$\equiv\sum_{l}M_{z}^{(l)},$
where $\rm o/e$ for occupied/empty states, $|u_{n\textbf{k}}\rangle$ is the periodic part of $|\psi_{n\textbf{k}}\rangle$ and $\epsilon_{z\alpha\beta}$ with $\alpha,\beta=x,y$ is the antisymmetric tensor. The layer-resolved magnetization of the $l$-th layer $M_{z}^{(l)}$ is then defined by summing over all the degrees of freedom except the layer index. For illustration, we model the 2D Chern-layer constituents of 3D Chern insulator by 2D Haldane model. Without loss of generality, the Chern number for each layer is set to be $C=1$ and interlayer coupling is weak such that the bulk gap remains finite. Details of parameters is shown in Fig. 1(b). In our numerical calculations, we consider two types of stacking, $AA$ and $AB$, as illustrated in Fig. 1(c) and the slab consists of five or six layers. The k mesh is set to $100\times 100$. The chemical potential $\mu$ is determined by looking for charge neutrality point in the same system with open boundary conditions. Then, $\alpha_{zz}^{(l)}$ can be obtained by numerically taking the ratio between change of $M_{z}^{(l)}$ and $\mathcal{E}_{z}$. We find that $M_{z}^{(l)}$ varies linearly for weak electric field for both types of stacking (see Fig. 1 in Supplemental Material [31]). The obtained $\alpha_{zz}^{(l)}$ is quantized to integers if $N_{\rm L}$ odd or to half-integers if $N_{\rm L}$ even, as shown in Table 2. The distribution is symmetric to the central plane of the system and the total $\alpha_{zz}$ is indeed vanishing as dictated by mirror symmetry. The quantization is limited to $1\%$ precision due to the finite-size effect [31].

### v3

@@ Abstract
Orbital magnetoelectric effect is closely related to the band topology of bulk crystalline insulators. Typical examples include the half quantized Chern-Simons orbital magnetoelectric coupling in three dimensional (3D) axion insulators and topological insulators, which are the hallmarks of their nontrivial bulk band topology. While the Chern-Simons coupling is well defined only for insulators with zero Chern number, the orbital magnetoelectric effects in 3D Chern insulators with nonzero (layer) Chern numbers are still open questions. In this work, we propose a never-mentioned quantization rule for the orbital magnetoelectric response in 3D Chern insulators, the spatial gradient of which is exactly quantized in unit of $e^{2}/h$. By theoretical analysis and numerical simulations, we demonstrate that such quantized orbital magnetoelectric response is exact for various types of interlayer hoppings and stackings, and remains robust even against disorder and lack of crystalline symmetries. We argue that the exact quantization has a topological origin and is protected by Chern number. Furthermore, we propose two promising material platforms to observe the proposed quantized orbital magnetoelectric response thanks to recent experimental developments in detecting spatial magnetic-field distributions in device systems.
I Introduction
@@
Band insulators can be classified into different topological phases according to different topological invariants [1, 2, 3, 4]. One important and basic topological invariant is the Chern number for two dimensional (2D) systems [2, 5, 1], which describes the winding number of hybrid Wannier centers within 2D Brillouin zone [6, 7]. Those time-reversal-broken band insulators with nonzero Chern number are called Chern insulators, which would exhibit quantum anomalous Hall effect [8, 9, 10]. The notion of 2D Chern insulators can be generalized to its 3D counterpart by stacking an infinite number of 2D Chern insulators along the vertical axis [11, 12, 13, 14, 15]. In principle, one can stack 2D Chern insulators along any direction so that one can define three independent Chern numbers $(C_{x},C_{y},C_{z})$, called Chern vector [16, 17], to characterize 3D band insulators. If the interlayer hopping is weak enough such that the bulk gap remains opened, such a topologically non-trivial 3D band insulator, named 3D Chern insulator, is adiabatically connected to an infinite number of decoupled 2D Chern-insulator layers.
@@
Topological 3D band insulators are known to exhibit non-trivial orbital magnetoelectric (OME) responses. The OME coupling $\alpha_{\mu\nu}$ is defined as the change in bulk electric polarization $P_{\mu}$ while applying magnetic field $B_{\nu}$, or equivalently the variation of bulk orbital magnetization $M_{\nu}$ induced by external electric field $\mathcal{E}_{\mu}$: $\alpha_{\mu\nu}=\partial P_{\mu}/\partial B_{\nu}=\partial M_{\nu}/\partial\mathcal{E}_{\mu}$ with $\mu,\nu=x,y,z$. This response is odd under inversion and time-reversal symmetry so that it vanishes if either of two symmetries is present. 3D time-reversal-protected topological insulators are classified by $\mathbb{Z}_{2}$ invariant [4, 18, 19, 20], which exhibit a half quantized bulk diagonal Chern-Simons OME response $\alpha_{\mu\nu}=e^{2}/2h\,\delta_{\mu\nu}$, modulo an integer multiple of $e^{2}/h$, once the surface state is gapped out [21, 22]. However, the Chern-Simons coupling is well defined only for bulk band insulators with zero Chern number within any 2D plane in momentum space, namely with vanishing Chern vector. This is because the expression of the Chern-Simons OME coupling coefficient involves an integration of Chern-Simons 3-form defined by the occupied Bloch states over the 3D Brillouin zone [22, 21, 23, 24, 25], which requires the existence of a smooth and periodic gauge throughout the 3D Brillouin zone [26, 27, 28]. However, such a gauge is non-existing for 3D Chern insulators due to topological obstruction.
@@
Nevertheless, nothing forbids 3D Chern insulators to have a topological OME response. Intuitively, let us consider a 3D Chern insulator with Chern vector $\textbf{C}=(0,0,C_{z})$, where $C_{z}=C(C\in\mathbb{Z},C\neq 0)$ with interlayer distance $d$. On the one hand, it is straightforward to see that the quantized anomalous Hall current in response to weak in-plane electric field would induce an off-diagonal OME response with its coefficient varying linearly in real space (see Sec. VII in Supplemental Materials [29]). On the other hand, if an out-of-plane uniform electric field ($\mathcal{E}_{z}$) is applied, the surface-state occupation would be different for every layer due to the vertical electrostatic potential drop. Then, the uneven distribution of electrons among the chiral edge states of each weakly coupled Chern-insulator layer would cause a layer-resolved orbital magnetization ($M_{z}$), which may manifest as a new type of diagonal OME effect unique to 3D Chern insulators, as schematically shown in Fig. 1(a). In this work, we find a never-mentioned quantization rule existing for the layer-resolved OME coupling of 3D Chern insulators: the difference between layer-resolved OME response coefficients is exactly quantized in units of $-Ce^{2}/h$ (see Eq. (7) and (8)). Such quantization rule is exact in the presence of various types of interlayer couplings and stacking ways, and even remains robust against breaking of crystalline symmetries and disorder. Furthermore, through extensive first principles density functional theory (DFT) calculations, we suggest Mn(Bi/Sb)2Te4[(Bi/Sb)2Te3]n ($n=0,1,2,...$) and Cr-doped (Bi/Sb)2Te3 as two classes of promising material platform (see Fig. 1(c) for schematic illustration) to realize such quantized magnetoelectric response. We further propose a specific experimental workflow using state-of-the-art magnetometer techniques to reveal the quantization rule in quasi-3D Chern insulator slabs.
Figure 1: (a) Schematic illustration for the origin of the quantization rule Eq. (8) for multilayer-layer Chern insulator slab: an out-of-plane electric field induces linearly increasing layer-projected out-of-plane orbital magnetization by redistributing electrons among the chiral edge states of different layers. (b) Honeycomb lattice for 2D Haldane model on which the hopping $t_{1}$ and $t_{2}\exp(i\phi)$ are marked. We study two ways of stacking: $AA$-stacking and $AB$-stacking. Unless mentioned ad hoc, we always use $t_{1}=1$, $t_{2}=1/3$, $\phi=\pi/3$, $t_{3}=0.2$, $t_{4}=0.15$ and the staggered mass term is zero. (c) Experimental proposal to observe the quantized OME response in a device formed by alternately stacking spacer and three-layer MnBi2Te4 (3L MBT) Chern layers, whose lattice structure is also given in the right panel, using SQUID-on-tip technique [30].
II Layer-resolved orbital magnetization
II.1 Definitions of layer-resolved orbital magnetization and layer Chern number
@@
Although the total diagonal $\alpha_{zz}$, only relevant in 3D case, is vanishing in the presence of mirror or inversion symmetry, the layer-resolved $\alpha_{zz}^{(l)}$ (with layer index $l$), which measures the change of layer-resolved orbital magnetization $M_{z}^{(l)}$ under a weak $\mathcal{E}_{z}$, may be finite. We first show the quantization rule for $\alpha_{zz}^{(l)}$ using the definition $\partial M_{z}^{(l)}/\partial\mathcal{E}_{z}$. Since $\mathcal{E}_{z}$ would break the translational symmetry in the $z$ direction, it is natural to consider a slab geometry. Then, the layer-resolved magnetization of the $l$-th layer $M_{z}^{(l)}$ can be defined by summing over all the quantum degrees of freedom except the layer index , which is a straightforward extension from the local definition of orbital magnetization [31, 32].
Specifically, we use two expressions for layer-resolved orbital magnetization in a slab. More details on their derivation can be found in Sec. I  of Supplemental Materials [29]. The first expression is derived from classical electrodynamics:
$M_{z}=\frac{-e}{2A}\sum_{E_{n}<\mu}\langle\phi_{n}|\textbf{r}\times\textbf{v}|\phi_{n}\rangle\bigg{|}_{z}\,,$ (1)
where r is the position, v is the velocity, $A$ is the total area, $e>0$ is the elementary charge, and $|\phi_{n}\rangle$s are the eigenstates with energy $E_{n}$ below chemical potential $\mu$. Note that this expression only applies to systems with open boundary condition (OBC). This is because position operators are ill-defined in the thermodynamic limit and terms like $\langle\phi_{n}|\textbf{r}|\phi_{n}\rangle$ diverge. This formula is particularly useful when we introduce disorder in the system in OBC. It contains both the contribution of circulation of the current density from the bulk and the contribution from the surface. Then, one can define layer-resolved orbital magnetization as
where the atomic lattice basis $|sl,\textbf{R}\rangle$ represents the Wannier basis state at layer $l$, sublattice $s$ of unit-cell R. Here, $P$ is the projection operator onto the occupied subspace.
In the thermodynamic limit with periodic boundary conditions, we need to define a layer-projected orbital magnetization $M_{z}^{(l)}$ using the expression of total orbital magnetization [32, 31]
$M_{z}=\frac{e}{\hbar A}{\rm Im\,Tr}\{PxQ(H-\mu)QyP-QxP(H-\mu)PyQ\},$ (3)
@@
where $Q=1-P$ is the projection operator onto the unoccupied subspace. The chemical potential $\mu$ is at charge neutrality and $x,y$ are 2D position operators. “Tr” in the above equation means taking the trace over the whole Hilbert space. Then, we can express $M_{z}$ in the Bloch eigen-function basis
@@
where $s$ means sublattice index in the unit cell, $l$ layer index, $N_{\rm L}$ number of 2D layers in the $z$ direction, $n$ band index with energy $E_{n\textbf{k}}$ at wavevector k, and $|sl,\mathbf{k}\rangle$ is the Fourier transform of the Wannier basis states $|sl,\textbf{R}\rangle$. In this basis, we can define a layer-projected orbital magnetization $M_{z}^{(l)}$
@@
$M_{z}^{(l)}=$ $\frac{e}{2\hbar A}{\rm Im}\big{\{}\epsilon_{z\alpha\beta}\sum_{s}\sum_{\textbf{k}}\sum_{\begin{subarray}{c}n,n^{\prime}\in{\rm o},\\ m\in{\rm e}\end{subarray}}C_{sl,n}(\textbf{k})C_{sl,n^{\prime}}^{*}(\textbf{k})\times(E_{m\textbf{k}}-\mu)\langle u_{n\textbf{k}}|\partial_{k_{\alpha}}u_{m\textbf{k}}\rangle\langle\partial_{k_{\beta}}u_{m\textbf{k}}|u_{n^{\prime}\textbf{k}}\rangle\big{\}}$ (5)
@@
$-\frac{e}{2\hbar A}{\rm Im}\big{\{}\epsilon_{z\alpha\beta}\sum_{s}\sum_{\textbf{k}}\sum_{\begin{subarray}{c}m,m^{\prime}\in{\rm e},\\ n\in{\rm o}\end{subarray}}C_{sl,m}(\textbf{k})C_{sl,m^{\prime}}^{*}(\textbf{k})\times(E_{n\textbf{k}}-\mu)\langle u_{m\textbf{k}}|\partial_{k_{\alpha}}u_{n\textbf{k}}\rangle\langle\partial_{k_{\beta}}u_{n\textbf{k}}|u_{m^{\prime}\textbf{k}}\rangle\big{\}}\,,$
where $A$ is the system area, $\rm o/e$ for occupied/empty states, $|u_{n\textbf{k}}\rangle$ is the periodic part of the Bloch eigenstate $|\psi_{n\textbf{k}}\rangle$ and $\epsilon_{z\alpha\beta}$ with $\alpha,\beta=x,y$ is the antisymmetric tensor.
In the same spirit, we can also define layer Chern number in the Bloch basis
@@
$=-\frac{\pi}{A}\epsilon_{z\alpha\beta}\,{\rm Im}\left\{\sum_{s\,l}\sum_{n,n^{\prime}\in{\rm occ}}\sum_{m\in{\rm emp}}\sum_{\textbf{k}}C_{sl,n}(\textbf{k})C_{sl,n^{\prime}}^{*}(\textbf{k})\langle u_{n\textbf{k}}|\partial_{k_{\alpha}}u_{m\textbf{k}}\rangle\langle\partial_{k_{\beta}}u_{m\textbf{k}}|u_{n^{\prime}\textbf{k}}\rangle\right.$
$\left.\quad\quad-\sum_{s\,l}\sum_{n\in{\rm occ}}\sum_{m,m^{\prime}\in{\rm emp}}\sum_{\textbf{k}}C_{sl,m}(\textbf{k})C_{sl,m^{\prime}}^{*}(\textbf{k})\langle u_{m\textbf{k}}|\partial_{k_{\alpha}}u_{n\textbf{k}}\rangle\langle\partial_{k_{\beta}}u_{n\textbf{k}}|u_{m^{\prime}\textbf{k}}\rangle\right\}$
$\equiv\sum_{l}C_{z}(l).$
Similar definitions are also used in Ref. [33, 34, 35].
II.2 Response of layer-resolved orbital magnetization to vertical electric field
For illustration, we model the 2D Chern-layer constituents of 3D Chern insulator by 2D Haldane model. Without loss of generality, the Chern number for each layer is set to be $C=1$ and interlayer coupling is weak such that the bulk gap remains finite. Details of parameters are shown in Fig. 1(b) and we consider two types of stacking, $AA$ and $AB$, as illustrated in Fig. 1(b). The k mesh is set to $100\times 100$. The chemical potential $\mu$ is determined by looking for charge neutrality point in the same system with open boundary conditions.
In a slab system, we calculate $\alpha_{zz}^{(l)}$ using $\partial M_{z}^{(l)}/\partial\mathcal{E}_{z}$. We employ two approaches to calculate $\alpha_{zz}^{(l)}=\partial M_{z}^{(l)}/\partial\mathcal{E}_{z}$ in a slab system. The first one is to calculate $M_{z}^{(l)}$ in the absence and in the presence of weak electric field using Eq. (5), then get $\alpha_{zz}^{(l)}$ by finite difference. In practice, we calculate $M_{z}^{(l)}$ for a series of weak electric field such that the response is still in the linear regime. Then, we fit the results to get the slope of the straight line and thus $\alpha_{zz}^{(l)}$. We use this approach for a slab using Eq. (5), in which the wavefunctions and eigenergies are calculated using periodic boundary condition (PBC) in the $x,y$ plane. But, the chemical potential appearing in Eq. (5) is still determined using a slab with open boundary condition (OBC) in all three spatial directions with $100\times 100$ sites in the $(x,y)$ plane due to the presence of gapless surface states.
The second approach is to calculate $\alpha_{zz}^{(l)}$ directly applying perturbation theory to Eq. (5). Specifically, when a vertical electric field is applied to a slab, our perturbed Hamiltonian reads: $H=H_{0}+e\mathcal{E}_{z}z$. Suppose the electric field is so weak such that bulk gap is not closed and the potential drop between the top and bottom layer is the smallest energy scale in the system. To get a perturbation theory of layer-resolved OME response, we need to express Eq. (5) of perturbed Hamiltonian $H$ using the information of $H_{0}$. By singling out the terms proportional to $\mathcal{E}_{z}$, the layer-resolved OME response $\alpha_{zz}^{(l)}$ is then derived by $\partial M_{z}^{(l)}/\partial\mathcal{E}_{z}$. After gathering all the first-order corrections (including energy, wavefunction and chemical potential with particle number fixed), a semi-analytic formula for $\alpha_{zz}^{(l)}$ can be derived. Details of the derivation are given in Sec. III  of Supplemental Materials [29]. We say “semi-analytic” since the determination of the chemical potential still requires solving numerically the system with open boundary condition, which is unavoidable due to the presence of gapless surface states in 3D Chern insulators.
Compared to the first approach, one advantage of the perturbative approach is that the quantization gradually improves as the $\mathbf{k}$ mesh becomes finer, as will be presented in the following. Another advantage of the perturbation approach is that the mirror symmetry is enforced in the results: $\alpha_{zz}^{(l)}$ is rigorously symmetric with respect to the central plane.
We apply the above two numerical approaches to multilayer Haldane model with both $AA$ and $AB$ stackings. It turns out that $M_{z}^{(l)}$ varies linearly for weak electric field for both types of stacking (see Fig. S1 in Supplemental Materials [29]). The obtained $\alpha_{zz}^{(l)}$ is quantized to integers if the number of layers $N_{\rm L}$ is odd or to half-integers if $N_{\rm L}$ is even, as shown in Table 1. The distribution is symmetric to the central plane of the system and the total $\alpha_{zz}$ is indeed vanishing as dictated by mirror symmetry. The quantization is limited to $1\%$ precision due to the finite-size effect [29]. If $\mathcal{E}_{z}$ is so weak that the electrostatic potential drop between the top and bottom layers can be treated as a perturbation to the system, then $\alpha_{zz}^{(l)}$ can be determined semi-analytically using perturbation theory as described above. Therefore, one can directly calculate $\alpha_{zz}^{(l)}$ using only the energy and Bloch functions of the unperturbed system. As shown in the parenthesis of Table 1, $\alpha_{zz}^{(l)}$ obtained by semi-analytic perturbation theory follow exactly the same quantization rule as those calculated by numerical finite-difference method.